OpenStudy (anonymous):

if h(t) represents the height of an object above ground level at time t and h(t) is given by h(t)=-16t^2+14t+1 find the height of the object at the time when the speed is zero.

OpenStudy (anonymous):

Hint: speed is zero=means when it is not moving.

OpenStudy (whpalmer4):

One trivial solution is h = 0 :-)

OpenStudy (anonymous):

Differentiating with respect to,we will get speed. Given that speed is zero. Then we can find time when speed.Equating the gotteb in the equation we will get height.

OpenStudy (anonymous):

\[h(t)=-16t ^{2}+14t+1\]

OpenStudy (anonymous):

\[d[h(t)] \div dt=-(16*2)t+14\]

OpenStudy (anonymous):

Here d[h(t)]÷dt=0

OpenStudy (anonymous):

Therefore, 0=-32t+14 t=14/32 Equating t=14/32 in given equation h(t)=−16t^2+14t+1 h(t)=\[16*(14/32)^{2}+14(14/32)+1\] =10.1875